### v1

# Deconstructing Obfuscation: A four-dimensional framework for evaluating Large Language Models assembly code deobfuscation capabilities

Abstract
Large language models (LLMs) have demonstrated remarkable capabilities across various software engineering tasks, yet their effectiveness for security-critical applications like binary analysis remains largely unexplored. This paper presents the first comprehensive evaluation of state-of-the-art (as of March 2025) commercial LLMs for assembly-level code deobfuscation, a task traditionally requiring specialized expertise and tools. We systematically tested seven models against four obfuscation scenarios: three individual techniques (bogus control flow, instruction substitution, and control flow flattening) and their combination. Our evaluation reveals striking performance variations across techniques and models, from autonomous deobfuscation to complete failure, that cannot be explained by model size or architecture alone. We propose a novel theoretical framework based on four distinct yet interconnected dimensions - Reasoning Depth, Pattern Recognition, Noise Filtering, and Context Integration — that provides explanatory power for these variations. Our analysis identifies five distinct error patterns that appear consistently across models: predicate misinterpretation, structural mapping errors, control flow misinterpretation, arithmetic transformation errors, and constant propagation errors. These patterns reveal fundamental limitations in how current LLMs process obfuscated code. Based on empirical findings, we establish a three-tier resistance model categorizing obfuscation techniques according to their effectiveness against LLM-based analysis: bogus control flow (low resistance), control flow flattening (moderate resistance), and instruction substitution or combined techniques (high resistance). This classification offers actionable guidance for both attack and defense scenarios. The universal failure against combined techniques defines a clear upper bound on current capabilities, demonstrating that sophisticated protection mechanisms remain effective against even advanced LLMs. Our framework creates a foundation for evaluating emerging capabilities, developing more resistant obfuscation techniques, and understanding how LLMs might transform security workflows. Rather than complete automation, our findings point to a new paradigm of human-AI collaboration in which LLMs reduce expertise barriers for certain aspects of reverse engineering while still requiring human guidance for complex deobfuscation tasks.
Keywords Large Language Models, Binary Deobfuscation, Dimensional Framework, Assembly Analysis, AI-assisted Security, Reverse Engineering
1 Introduction
Code obfuscation involves the transformation of a program into a form intentionally more difficult to understand, targeting both human analysts and automated analysis tools [20]. While the functionality remains identical, the syntax and structure are deliberately altered to obscure the code’s true purpose. These transformations range from simple techniques, such as renaming identifiers or inserting meaningless instructions, to advanced methods, such as control flow flattening or full program encryption [21]. Obfuscation serves legitimate applications, such as protecting intellectual property by preventing reverse engineering, but it is often exploited by malware creators to hide malicious functionality, complicating detection and analysis [22]. This dual use scenario, which serves both defensive software protection and offensive malware development, positions obfuscation as a critical area of cybersecurity research [29].
Reverse engineering binary code typically begins with decompilation - the process of transforming machine code back into higher-level source code representations [30]. Decompilation itself is a complex challenge, requiring the reconstruction of control flow, data types, and program structure from binaries where variable names, comments, and logical organization have been lost during compilation [31]. Building upon decompilation, traditional deobfuscation methods focus on reversing or simplifying the obfuscation to clarify the underlying logic [24]. These typically involve program analysis techniques: static (analyzing code without executing it) [23] and dynamic (running or emulating code) [25, 26]. Static methods employ tools such as IDA [27] or Ghidra [28], analyzing control flow to recognize high-level constructs. Advanced static analysis techniques such as symbolic execution, abstract interpretation, or program slicing analyze logic without execution. Dynamic analysis methods include controlled execution environments or taint analysis to reveal hidden logic at runtime. Usually, hybrid approaches that combine static and dynamic analysis offer the best results. However, traditional methods often require significant manual intervention and struggle against advanced obfuscation, which can render common analysis tools ineffective. This limitation results in an ongoing automation challenge, an ’arms race’ between code protectors and analysts [22].
Large language models (LLMs) have made substantial advances in recent years, transforming natural language processing through models such as BERT [1], GPT [2], and T5 [3]. Using transformer architectures [4], extensive datasets such as The Stack [16], and enhanced computational resources, LLMs now excel in various complex tasks across multiple domains. In software engineering specifically, models like CodeBERT [5] and Codex [17] — the foundational technology behind GitHub Copilot — have demonstrated exceptional performance in code completion, error detection, and multilingual code generation. Recent comprehensive surveys [15] and empirical analyses [17] highlight both their significant capabilities and ongoing limitations. Techniques such as instruction tuning [6] and in-context learning [2] have further boosted their effectiveness. Furthermore, rigorous benchmarks such as RepoBench [7] and LiveCodeBench [8] systematically evaluate their performance, while research on attention mechanisms [9] and model robustness [10] underscores their potential for secure coding applications. Emerging tools such as CodeGen [11] and CodeT5+ [19, 12], alongside ongoing studies [13, 14], continue to push the limits of automated code manipulation, highlighting the intrinsic connection between advancements in natural language processing and programming language analysis [18].
These advancements in code manipulations using large language models (LLMs) might introduce new opportunities for automating deobfuscation tasks. Imagining deobfuscation as a translation task - similar to translating between natural languages - these models can produce more intelligible forms from obfuscated code. Previous studies have successfully used neural machine translation (seq2seq) models to decompile assembly back into high-level languages like C or Go [32], achieving accuracy comparable to traditional decompilers. Such models can rapidly adapt across multiple languages, demonstrating substantial potential for automated reverse engineering.
Early explorations into LLM-driven deobfuscation show promise. Lachaux et al. introduced DOBF, a pre-training method that helps models recover source code from obfuscated snippets [33]. Although still limited, DOBF underscores the potential for significant future advances. Unlike deterministic deobfuscators, LLMs provide semantic summaries and high-level interpretations of obfuscated logic, significantly aiding human analysts. However, current LLMs remain prone to inaccuracies or "hallucinated" logic, risking misleading interpretations. However, combining these models with traditional tools has already shown substantial benefits, significantly reducing manual analysis effort [34].
Our research specifically addresses this gap by systematically evaluating current state-of-the-art(as of March 2025) general-purpose LLMs on raw assembly code from real-world obfuscated binaries. Although promising, most previous studies on LLM-driven deobfuscation have focused primarily on high-level languages or direct binary-to-source decompilation, often relying heavily on domain-specific fine-tuning and specialized knowledge such as syntax trees or predefined language constructs [35]. This highlights the need to evaluate general-purpose LLMs directly on disassembled assembly code, a common, yet challenging scenario frequently encountered during binary analysis. We specifically target disassembled native assembly code, which analysts frequently face when high-level decompilers fail. By treating raw assembly instructions as token sequences—analogous to natural language sentences—we uniquely test whether large-scale general-purpose LLMs, trained on broad text and code datasets, can practically automate assembly-level deobfuscation without fine-tuning or specialized domain knowledge, significantly reducing manual analysis efforts. This perspective provides novel insights bridging high-level deobfuscation and detailed binary analysis, highlighting capabilities and limitations previously unidentified.
Based on our analysis of current capabilities and gaps in LLM-driven deobfuscation, this research specifically addresses the following questions:
  • To what extent can general-purpose LLMs successfully analyze and deobfuscate raw assembly code from obfuscated binaries without specialized fine-tuning, and what levels of human intervention are required for different models and techniques?
  • How do different obfuscation techniques (bogus control flow, instruction substitution, control flow flattening, and combined techniques) affect LLM deobfuscation performance, and what dimensional capabilities (reasoning depth, pattern recognition, noise filtering, context integration) do they primarily challenge?
  • What specific types of error do LLMs make when attempting to deobfuscate assembly code, and how do these error patterns relate to fundamental limitations in LLMs’ reasoning and pattern recognition capabilities?
  • What implications do current LLM deobfuscation capabilities have for developing next-generation obfuscation techniques and improved automated deobfuscation tools in the cybersecurity landscape?
2 Methodology
We evaluated eight state-of-the-art commercial Large Language Models (LLM) - GPT-3o Mini High, GPT-4o, GPT-4.5, O1 Pro Mode, DeepSeekR1, Grok3, Grok2, and Claude 3.7 Sonnet - on their ability to analyze and deobfuscate a known C program obfuscated using Obfuscator-LLVM (OLLVM) [37]. These models differ significantly in several dimensions, such as their model size, the scale and nature of their training data sets, the lengths of the context window, and the inherent reasoning capabilities. A detailed comparison of the specific characteristics of each model is provided in Table 1.
Table 1: Summary of LLM Properties and Reasoning Mechanisms
Model CW MOT BR
GPT-4o [41] 128,000 16,384 No
GPT-3o Mini High [41] 200,000 100,000 Yes
GPT-4.5 Preview [41] 128,000 16,384 No
o1 [41] 200,000 100,000 Yes
Claude 3.7 Sonnet [40] 200,000
Normal: 8,192
Extended: 64,000
Yes
DeepSeek R1 [45] 131,072 131,072 Yes
GROK 3 [42] 1,000,000 128,000 Yes
GROK 2 [43] 131,072 28,000 Yes
  • Legend: Context Window = CW, Max Output Tokens = MOT, Built-in Reasoning = BR. DeepSeek R1
This specific test case was previously comprehensively analyzed by Quarkslab [36] and was chosen due to its documented complexity and practical relevance, making it an ideal benchmark for a semi-realistic evaluation scenario. Quarkslab successfully reversed OLLVM’s obfuscation protections, such as control-flow flattening, bogus control-flow, and instruction substitution, by employing symbolic execution within the Miasm framework, though their approach required considerable manual effort and specialized expertise to analyze the obfuscated code. Furthermore, the chosen test program includes numerous conditional branches, making it particularly suitable for assessing the effectiveness and resilience of different obfuscation methods.
The original unobfuscated C function used in our evaluation is shown in Listing 1. This function implements a straightforward algorithm that computes different arithmetic and bitwise transformations based on the input value modulo 4. Despite its relative simplicity, the function incorporates several characteristics that make it ideally suited for obfuscation testing: multiple conditional branches, diverse bitwise operations, and a distinctive magic constant (0xBAAAD0BF) that serves as a recognizable marker throughout the assembly code.
Listing 1: Original unobfuscated C function
1unsigned int target_function(unsigned int n) {
2 unsigned int mod = n % 4;
3 unsigned int result = 0;
4
5 if (mod == 0)
6 result = (n | 0xBAAAD0BF) * (2 ^ n);
7 else if (mod == 1)
8 result = (n & 0xBAAAD0BF) * (3 + n);
9 else if (mod == 2)
10 result = (n ^ 0xBAAAD0BF) * (4 | n);
11 else
12 result = (n + 0xBAAAD0BF) * (5 & n);
13
14 return result;
15}
We compiled this function using OLLVM, applying various obfuscation configurations to produce five distinct binaries. Each binary was disassembled using Capstone [38], producing x86_64 assembly code for our detailed low-level analysis. For reproducibility of our results, we provide the specific compilation flags used for each binary:
  • code_unobf.capstone: Baseline with no obfuscation.
  • code_sub.capstone: Instruction substitution applied. Compiled with the following flags: -mllvm -sub
  • code_fla.capstone: Control flow flattening applied. Compiled with flags: -mllvm -fla -mllvm -perFLA=100
  • code_bcf.capstone: Bogus control flow applied. Compiled with flags: -mllvm -bcf -mllvm -boguscf-prob=100 -mllvm -boguscf-loop=1
  • code_all.capstone: All three obfuscation techniques combined. Compiled with flags: -mllvm -sub -mllvm -fla -mllvm -perFLA=100 -mllvm -bcf -mllvm -boguscf-prob=100 -mllvm -boguscf-loop=1
We specifically focused on x86_64 architecture as it remains the dominant instruction set in desktop and server environments, making it particularly relevant for real-world malware analysis and reverse engineering scenarios.
Our evaluation specifically targeted widely available commercial LLMs due to their advanced capabilities and relevance to practical cybersecurity scenarios. Initially, our goal was to perform a statistical analysis to quantify each model’s effectiveness in deobfuscation tasks. Our methodical testing established that certain obfuscation methods were entirely resistant to deobfuscation by these models. We implemented a comprehensive qualitative analysis to thoroughly document the specific errors exhibited by each model. To ensure robustness, each model and scenario combination was extensively tested multiple times to identify common patterns in both successful and failed decompilation attempts. This systematic approach allowed us to select the most representative results for a detailed review. The repeated testing methodology not only improved the reliability of our findings, but also closely simulated realistic attacker constraints, including limited knowledge, resources, and opportunities for experimentation.
We structured interactions with the AI models based on incremental attacker knowledge levels defined as follows:
  • Level 0: No Knowledge Needed—AI fully deobfuscates without assistance.
  • Level 1: Basic Guidance—Minimal hints to correct minor errors.
  • Level 2: Structural Correction—Significant guidance needed for structural issues.
  • Level 3: Major Intervention—Detailed guidance necessary to resolve complex logic errors.
  • Level 4: Expert Rework—Extensive expert intervention required.
  • Level 5: Beyond Expert Correction—Errors too fundamental, requiring a complete restart.
  • No Level(-): Unable to Analyze—AI produced no meaningful output.
Detailed dialogues with each model were documented and analyzed to identify patterns in reasoning, common mistakes, and overall effectiveness in the deobfuscation process.
Ultimately, our primary objective is to evaluate how commercial AI models might realistically be exploited for code analysis and deobfuscation. This has dual implications for cybersecurity: it highlights risks to legitimate software protection mechanisms while also informing potential defensive applications in malware analysis. These findings aim to help legitimate software developers create more resilient obfuscation techniques and security researchers understand the evolving landscape of AI-assisted code analysis tools.
3 Results
The performance of the models evaluated against different obfuscation methods is summarized in Table 2. Each entry indicates the level of attacker knowledge required for successful deobfuscation, based on the criteria defined in Section 2. Higher numerical values indicate greater difficulty and increased human intervention. Also, "-" indicates that the particular scenario is was not applicable.
Table 2: Obfuscation Variants and Required Attacker Knowledge Levels
Model BCF IS CFF All
GPT-4o 4 4 1 -
GPT-4.5 1-2 4 0 5
GPT-Pro-o1 3 - 0 5
GPT-3o Mini 4-5 - - -
DeepSeekR1 5 5 1-2 5
Grok3 1 4 0 5
Grok2 - - 1-2 -
Claude 3.7 Sonnet 0 1-2 0-1 5
  • Legend: Bogus Control Flow = BCF, Control Flow Flattening = CFF, Instruction Substitution = IS.
The following sections present a qualitative analysis of model performance against four obfuscation techniques: Bogus Control Flow (BCF), Instruction Substitution (IS), Control Flow Flattening (CFF), and Combined Techniques (All). Our analysis uses the comprehensive Quarkslab analysis [36] as an implicit baseline—their detailed technical breakdown successfully deobfuscated the same code through symbolic execution, but required specialized expertise and significant time investment. We focus on identifying specific error patterns rather than quantitative success rates, using standardized prompts across all models as described in Section 2. Each subsection addresses one obfuscation method, evaluating individual model performance and excluding models that produced no meaningful results. The complete transcripts of our interactions with LLM are not included in this article due to length constraints, but have been made publicly available in a dedicated repository[60] to support reproducibility and further analysis. The subsequent Discussion section (Section 4) synthesizes these findings into a theoretical framework explaining LLM deobfuscation capabilities and their implications for cybersecurity.
3.1 Bogus control flow
This section examines Bogus Control Flow (BCF). This obfuscation aims at complicating reverse engineering by introducing misleading control flow structures in the form of opaque predicates. For a detailed exposition and further insight into this method, see [36, 44]. Among the evaluated models, only Claude 3.7 Sonnet successfully deobfuscated the BCF-protected code on its initial attempt without requiring additional guidance; consequently, it is excluded from the detailed comparative analysis presented herein.
3.1.1 ChatGPT-4o
In our examination of the deobfuscation dialogue with ChatGPT-4o, we identified a significant error in its deobfuscation attempt. The mistake involved misinterpreting an opaque predicate. The incorrect interpretation of the model resulted in errors that predicted the code control flow.
The assembly code includes the condition (var1 * (var1 - 1)) & 1 == 0, where var1 is loaded from memory at MEMORY[0x404000]. This predicate is always true because the product of two consecutive integers is always even, ensuring the least significant bit is 0. ChatGPT initially misinterpreted this predicate, stating that if this condition or another condition (var2 < 10) was true (expressed as flag1 || flag2), the program would exit to “END_PROGRAM”. Conversely, it assumed that if the condition failed, the program would jump to address 0x40169e, initially labeling this as an exit, then later correctly identifying it as an infinite loop. However, since flag1 (the opaque predicate) always evaluates to true, the combined condition flag1 || flag2 is always true, making the infinite loop at 0x40169e unreachable.
Relevant assembly snippet is shown in Listing  2 and here is explanation of key steps:
  • imul eax, edx computes var1 * (var1 - 1).
  • and eax, 1 checks the least significant bit. Since the product is always even, this is always 0.
  • sete dh sets dh to 1 because the result of eax == 0 is always true.
  • test dh, 1 and jne 0x40104f ensure the jump to 0x40104f is always executed, rendering the subsequent jump to 0x40169e (the infinite loop) unreachable.
Listing 2: Relevant Assembly Snippet
1mov eax, [0x404000] ; eax = var1
2mov edx, eax
3sub edx, 1 ; edx = var1-1
4imul eax, edx ; eax = var1*(var1-1)
5and eax, 1 ; eax = 0 (always even)
6cmp eax, 0
7sete dh ; dh = 1 always
8or dh, al ; dh unchanged
9test dh, 1
10jne 0x40104f ; always jumps
11jmp 0x40169e ; dead code
Here ChatGPT-4o incorrectly assumed that this condition could fail, producing overly complicated pseudocode with unnecessary conditional checks. As a result provided completely incorrect deobfuscated code Listing  3 :
Listing 3: Final output ChatGPT-4o
1function main() {
2 int var1 = MEMORY[0x404000]
3 int var2 = MEMORY[0x404004]
4
5 // Compute opaque predicates
6 bool flag1 = ((var1 * (var1 - 1)) & 1) == 0
7 bool flag2 = (var2 < 10)
8
9 // Fake loop trap, never executes
10 if (!flag1 && !flag2) {
11 while (true) {
12 // Fake infinite loop (never reached)
13 }
14 }
15
16 // Real execution
17 execute_main_logic()
18}
19
20function execute_main_logic() {
21 // Perform real program operations
22}
ChatGPT-4o ultimately failed to deobfuscate the assembly accurately due to its misinterpretation of the opaque predicate.
3.1.2 ChatGPT 4.5
The analysis performed with ChatGPT 4.5 began with the examination of memory accesses at addresses 0x404000 (var1) and 0x404004 (var2). The model identified the following recurring assembly pattern shown in Listing 4:
Listing 4: Relevant Assembly Snippet
1mov eax, [0x404000] ; var1
2mov edx, eax
3sub edx, 1 ; edx = var1 - 1
4imul eax, edx ; eax = var1*(var1-1)
5and eax, 1 ; Check if even
6cmp eax, 0
7sete dh ; dh = 1 if even
ChatGPT 4.5 correctly identified this code snippet as an opaque predicate: the condition (var1 * (var1 - 1)) & 1 == 0 is always true because the product of two consecutive integers is even, thus ensuring branches leading to the address 0x40169e are unreachable. Upon explicitly providing the conditions var1 = 0 and var2 = 0, the model appropriately recognized instructions like cmp ecx, 0xa (checking var2 < 10) as obfuscation noise, since they always evaluate to true under these conditions.
However, ChatGPT 4.5 initially misrepresented the arithmetic operations involved. For instance, it initially produced the following incorrect transformation for case 0 as shown in Listing 5:
Listing 5: Incorrect Transformation for Case 0
result = (input | 0xbaaad0bf) * ((input | 0xbaaad0bf) ^ 2)
This statement differs significantly from the original assembly instructions. The correct interpretation, based on the original assembly at address 0x401138, is shown in Listing 6:
Listing 6: Correct Transformation for Case 0
result = (input | 0xbaaad0bf) * (input ^ 2)
After requesting verification, ChatGPT 4.5 subsequently corrected its output, providing the following accurate representation aligned with the assembly in Listing 7:
Listing 7: Final output 4.5
1int obfuscated_func(int input) {
2 int result, tmp1, tmp2;
3 switch (input & 3) {
4 case 0: tmp1 = input | 0xbaaad0bf;
5 tmp2 = input ^ 2;
6 result = tmp1 * tmp2; break;
7 case 1: tmp1 = input & 0xbaaad0bf;
8 tmp2 = input + 3;
9 result = tmp1 * tmp2; break;
10 case 2: tmp1 = input ^ 0xbaaad0bf;
11 tmp2 = input | 4;
12 result = tmp1 * tmp2; break;
13 case 3: tmp1 = input + 0xbaaad0bf;
14 tmp2 = input & 5;
15 result = tmp1 * tmp2; break;
16 }
17 return result;
18}
This final result correctly matches the logic implemented at addresses 0x401138 (case 0), 0x40128a (case 1), 0x4013dc (case 2), and 0x401486 (case 3). ChatGPT 4.5 corrected its initial errors with guidance, but required human intervention for accurate results.
3.1.3 ChatGPT-pro-o1
The ChatGPT-pro-o1 began its analysis by identifying a recurring conditional pattern involving memory locations 0x404000 (G1) and 0x404004 (G2), along with arithmetic transformations applied to a local variable (VALUE). The following sections detail its performance.
Initially, ChatGPT-pro-o1 correctly recognized four arithmetic transformations from the obfuscated assembly code. These transformations matched the ground truth precisely, despite the presence of obfuscation techniques such as redundant increment/decrement instructions and stack manipulations. The model also successfully identified the opaque predicate ((G1 * (G1 - 1)) & 1) == 0 || G2 < 10 as always evaluating to true, based on the mathematical property that the product of two consecutive integers is always even. However, despite correctly recognizing this predicate initially, ChatGPT-pro-o1 treated it as part of an iterative logic rather than a consistently true condition.
The model encountered the same assembly pattern previously shown in Listing 4. This condition, observed multiple times (e.g., at 0x401009 and 0x40108c), consistently evaluates to true at 0x401044. However, ChatGPT-pro-o1 initially interpreted these repetitions as indicative of essential iterative logic rather than obfuscation redundancy.
Furthermore, instead of identifying the concise switch-like structure at address 0x401074, triggered by data & 3, ChatGPT-pro-o1 initially proposed a complex chain of conditional checks with repeated transformations. In contrast, the actual assembly implements a straightforward conditional structure based on data & 3 at 0x401074, selecting exactly one transformation (such as the one at 0x401138) and then terminating via a single return instruction at 0x40169d. ChatGPT-pro-o1’s initial interpretation diverged by incorrectly modeling the process as iterative. After further prompting and clarification, ChatGPT-pro-o1 corrected its interpretation and provided an accurate deobfuscated implementation that correctly aligns with the actual assembly logic, structured as a switch statement dependent on input & 3.
ChatGPT-pro-o1 excelled at recognizing arithmetic patterns in obfuscated code but initially misread redundant control flow as iterative logic, correcting it with guidance.
3.1.4 Model o3-mini-high
The o3-mini-high began by analyzing recurring conditional checks involving global variables located at memory addresses 0x404000 (param1 = var1) and 0x404004 (param2 = var2). Specifically, it identified the same assembly sequence previously analyzed for ChatGPT-4.5, presented in Listing 4. The model correctly recognized this as an opaque predicate, ((param1 - 1) * param1) & 1 == 0, which consistently evaluates to true due to the mathematical property that the product of two consecutive integers is always even. As a result, the branches to the address 0x40169e were correctly identified as unreachable.
Despite correctly identifying this opaque predicate, the model subsequently represented the control flow incorrectly as a looping state machine. Its initial interpretation is presented in the pseudocode snippet below Listing 8:
Listing 8: o3-mini initial try
1function obfuscated_function():
2 while True:
3 flag1 = (((param1 - 1) * param1) & 1 == 0) or (param2 < 10)
4 if not flag1: goto cleanup_and_return
5 local_value = get_some_local_value()
6 if (control_byte & 1) == 0:
7 if not flag2: goto cleanup_and_return
8 local_value = transform_variant_A(local_value)
9 else:
10 if not flag3: goto cleanup_and_return
11 local_value = transform_variant_B(local_value)
12 local_value = complex_transform(local_value)
13 if termination_condition_reached(): break
14 cleanup_and_return:
15 return
After providing explicit conditions (param1 = 0, param2 = 0) and highlighting hidden constants (e.g., 0xbaaad0bf, 2, 3, 4, 5), the model produced a more structured version. However, it continued to interpret the assembly as a loop-based state machine. Further prompts requesting clarification and restructuring resulted in the following representation, which remained incorrect Listing  9:
Listing 9: o3-mini final try
1function obfuscated_function():
2 state = initial_state()
3 control1 = initial_control_byte()
4 while not termination_condition(state):
5 if (control1 & 1) == 1:
6 state = (state OR 0xbaaad0bf) * (state XOR 2)
7 else:
8 state = (state XOR 0xbaaad0bf) * (state OR 4)
9 control2 = compute_secondary_control(state)
10 if (control2 & 1) == 1:
11 state = (state AND 0xbaaad0bf) * (state + 3)
12 else:
13 state = (state + 0xbaaad0bf) * (state AND 5)
14 control1 = update_control(state)
15 return state
This final output differs significantly from the actual assembly logic, which implements a single-pass switch determined by the expression data & 3 at address 0x401074. The correct behavior, confirmed by manual analysis, executes exactly one arithmetic transformation (for example, at address 0x401138 for case 0) and terminates with a single return instruction at 0x40169d. However, the model repeatedly interpreted the assembly as an iterative construct, a structure not present in the original code. Despite clarifications, ’o3-mini-high’ persistently misread the non-iterative assembly as a loop, accurately spotting transformations but missing the single-pass structure.
3.1.5 Grok 3
In Grok 3 initial attempt, Grok 3 interpreted the assembly code as containing a functional loop driven by conditions involving variables A and B, generating the pseudocode shown in Listing 10:
Listing 10: Grok 3 initial interpretation
while (A > 1 AND B < 10): // Switch-like logic based on x & 3
However, considering the explicitly provided conditions A = 0 and B = 0, the loop condition (A > 1 AND B < 10) evaluates to false, because: A > 1 is false (since 0 > 1 is false), B < 10 is true (since 0 < 10 is true). Therefore, the combined condition (false AND true) evaluates to false.
Thus, under these explicit conditions, the loop never executes, contradicting Grok 3’s initial interpretation. Upon receiving further guidance explicitly stating that [0x404000] = 0 and [0x404004] = 0, Grok 3 revised its analysis, removing the loop structure and accurately producing the pseudocode shown in Listing 11:
Listing 11: Grok 3 final deobfuscation
1function process(x):
2 state = x & 3
3 if state == 0:
4 result = (x | 0xbaaad0bf) * (x ^ 2)
5 else if state == 1:
6 result = (x & 0xbaaad0bf) * (x + 3)
7 else if state == 2:
8 result = (x ^ 0xbaaad0bf) * (x | 4)
9 else: // state == 3
10 result = (x + 0xbaaad0bf) * (x & 5)
11 return result
This final pseudocode correctly reflects the assembly logic, which involves performing exactly one arithmetic transformation based on the condition x & 3. Each conditional jump involving variables A and B consistently results in a single execution path rather than iterative looping. Grok 3 accurately deobfuscated the code with explicit guidance, though initial errors suggest condition clarity is crucial.
3.1.6 DeepSeek R1
Similar to some other models, DeepSeek R1 incorrectly encased the deobfuscated logic within a while loop Listing  12:
Listing 12: DeepSeek R1 initial error
1while True:
2 # Initial condition check
3 even_check = (var_A * (var_A - 1)) & 1 # Basically var_A % 2
4 cond_flag = (even_check == 0) and (var_B < 10)
5
6 if not cond_flag:
7 break
Despite prompts to reassess its deobfuscation, R1 persisted with a loop, simplifying it to a for loop with exactly 10 iterations:
Listing 13: DeepSeek R1 incorrect loop
1for _ in range(10): # Exactly 10 iterations
2 remainder = result % 4
Although the switch statement itself was correctly interpreted, the erroneous outer loop caused the transformation to be applied 10 times, repeatedly altering the result variable as shown in Listing 14:
Listing 14: DeepSeek R1 Last Attempt
1def transform(input_val):
2 result = input_val
3 for _ in range(10): # Exactly 10 iterations
4 remainder = result % 4
5 if remainder == 0:
6 result = (result | 0xBAAAD0BF) * (result ^ 2)
7 elif remainder == 1:
8 result = (result & 0xBAAAD0BF) * (result + 3)
9 elif remainder == 2:
10 result = (result ^ 0xBAAAD0BF) * (result | 4)
11 else:
12 result = (result + 0xBAAAD0BF) * (result & 5)
13 return result
DeepSeek R1’s misinterpretation of the bogus control flow as requiring a concrete loop structure (forcing exactly 10 iterations) demonstrates a fundamental failure in recognizing the non-iterative nature of the original algorithm.
3.2 Instruction Substitution
In this section, we analyze the obfuscation known as Instruction Substitution (IS). This type of obfuscation aims at substituting binary operators with equivalent but more complex sequences of instructions. For a comprehensive description and additional details regarding this obfuscation technique, refer to [36, 44]. As shown in Table 2, GPT-Pro-o1 and GPT-3o Mini failed completely, unable to provide any meaningful results.
3.2.1 ChatGPT 4o
The ChatGPT 4o correctly identified the underlying switch structure based on the condition input & 3. However, the arithmetic and bitwise transformations derived by ChatGPT 4o differed substantially from the logic explicitly implemented by the assembly blocks located at addresses 0x27, 0xb7, 0x113, and 0x1ac. The final deobfuscation produced by ChatGPT 4o is shown in Listing 15:
Listing 15: Instruction Substitution o4 final deobfuscation attempt
1unsigned int function(unsigned int input) {
2 unsigned int result = 0;
3 switch (input & 3) {
4 case 0:
5 result = (input ^ 0xe6c98769) * ((input & 2) | 2);
6 break;
7 case 1:
8 result = 0;
9 break;
10 case 2:
11 result = (input ^ 0x811d5b51) * ((input & 4) | 4);
12 break;
13 case 3:
14 result = (input + 0xbaaad0bf) * (input ^ 0xf4577a2a);
15 break;
16 }
17 return result;
18}
A detailed analysis shows substantial inaccuracies in ChatGPT 4o’s approach as seen in Listing 15. In the first case, the model incorrectly introduced a bitwise XOR operation involving an unrelated constant (0xe6c98769), deviating from the correct bitwise OR operation with constant 0xBAAAD0BF as represented in the original assembly. The second case was significantly simplified by ChatGPT 4o, resulting in a constant zero, whereas the actual assembly involves necessary bitwise AND operations followed by arithmetic multiplication, neither of which were correctly represented by the model. In the third case, ChatGPT 4o erroneously introduced a constant (0x811d5b51) and incorrectly replaced the intended bitwise OR-based multiplication with alternative operations. Finally, in the fourth case, although the addition of the constant 0xBAAAD0BF was accurately identified, the model incorrectly applied a bitwise XOR operation with the constant 0xf4577a2a instead of the correct bitwise AND operation specified by the original assembly. Overall ChatGPT 4o failed to successfully deobfuscate the code.
3.2.2 GPT 4.5
The ChatGPT 4.5 correctly identified the switch structure conditioned on input & 3. However, the arithmetic and bitwise transformations it proposed differ significantly from those explicitly represented in the original assembly. Specifically, the final output from ChatGPT 4.5 is shown in Listing 16:
Listing 16: Instruction Substitution 4.5 final deobfuscation attempt
1unsigned int function(unsigned int input) {
2 switch (input & 3) {
3 case 0:
4 return ((~input ^ 0xBAAAD0BF ^ 0xE6C98769) + 2) * (input | 2);
5 case 1:
6 return ((~input | 0xBAAAD0BF) & 0x8ABD1CD5) * (input - 3);
7 case 2:
8 return ((input ^ 0x0F603E35 ^ 0xBAAAD0BF) + 0x811D5B55)
9 * (input ^ 0x0F603E35 ^ 0xBAAAD0BF);
10 case 3:
11 return (-(input + 0xBAAAD0BF)) * ((~input | 5) & 0xF4577A2A);
12 }
13 return 0; // Unreachable
14}
An in-depth analysis of ChatGPT 4.5’s proposed implementation in Listing 16 shows systematic deviations from the original logic. In the first case (address 0x27), the model incorrectly introduced bitwise inversion ( input), extraneous XOR operations (with constants 0xE6C98769 and 0xBAAAD0BF), and an additional arithmetic offset (+ 2). These operations are not present in the original assembly logic, which correctly employs a simple bitwise OR operation combined with XOR-based multiplication. In the second case (address 0xb7), the assembly correctly implements a bitwise AND operation followed by arithmetic addition. However, ChatGPT 4.5 inaccurately replaced these operations with bitwise inversion, OR-based operations involving unrelated masks (0x8ABD1CD5), and an arithmetic subtraction (input - 3), significantly deviating from the actual code behavior. For the third case (address 0x113), the original assembly clearly specifies a straightforward XOR operation using the constant 0xBAAAD0BF, multiplied by a bitwise OR operation involving the input. Conversely, ChatGPT 4.5 introduced multiple redundant XOR operations (e.g., 0x0F603E35), extraneous constants (0x811D5B55), and unnecessarily complex arithmetic, deviating substantially from the correct logic. Lastly, in the fourth case (address 0x1ac), rather than accurately applying a simple arithmetic addition followed by a bitwise AND operation, the model proposed incorrect arithmetic negation, unnecessary bitwise inversion, and irrelevant OR-based operations with an unrelated constant (0xF4577A2A), misrepresenting the original code significantly.
ChatGPT 4.5 failed to accurately deobfuscate the assembly due to significant errors in arithmetic and bitwise transformations despite identifying the switch structure.
3.2.3 Grok 3
Grok 3 made multiple attempts to deobfuscate the provided assembly code, while its correctly identified the switch structure based on input & 3, its interpretation of computational logic significantly deviated from the intended functionality. Grok 3’s final deobfuscation attempt resulted in the implementation shown in Listing 17:
Listing 17: Instruction Substitution Grok3 final deobfuscation attempt
1unsigned int function(unsigned int input) {
2 switch (input & 3) {
3 case 0: // input & 3 == 0
4 result = (input ^ 0xbaaad0bf) & 0xe6c98769;
5 break;
6 case 1: // input & 3 == 1
7 result = (input & 0xbaaad0bf) * (3 + input);
8 break;
9 case 2: // input & 3 == 2
10 result = (input ^ 0xf603e35) & 0xbaaad0bf;
11 break;
12 case 3: // input & 3 == 3
13 result = (input + 0xbaaad0bf) * (input & 5);
14 break;
15 }
16 return result;
17}
Upon detailed verification, Grok 3’s results shown in Listing 17 showed partial accuracy. Specifically, in cases 1 and 3, Grok 3 precisely matched the intended assembly logic. For case 1, it correctly implemented the bitwise AND operation with 0xBAAAD0BF, followed by arithmetic addition (3 + input) and multiplication. Similarly, in case 3, the arithmetic addition with 0xBAAAD0BF followed by multiplication with the result of a bitwise AND operation (input & 5) was accurately represented. In contrast, significant deviations occurred in cases 0 and 2. In case 0, Grok 3 incorrectly simplified the original logic. The intended assembly explicitly involves a bitwise OR with 0xBAAAD0BF and multiplication using an XOR operation with the constant 2. Grok 3, however, mistakenly replaced these operations with a single bitwise XOR operation (input ^ 0xBAAAD0BF) followed by an incorrect bitwise AND with an unrelated constant (0xE6C98769), thus significantly altering the intended arithmetic logic. For case 2, the intended assembly logic involves a bitwise XOR operation with 0xBAAAD0BF followed by arithmetic multiplication involving a bitwise OR operation with a small constant. Instead, Grok 3 inaccurately applied a bitwise XOR with an extraneous constant (0x0F603E35), followed by a bitwise AND with 0xBAAAD0BF. These inaccuracies substantially differ from the intended computational steps present in the assembly. So, despite success in some cases, Grok 3 ultimately failed to deobfuscate the assembly due to critical errors in computational logic.
3.2.4 DeepSeek R1
The DeepSeek R1 correctly identified the high-level switch structure based on input & 3, but demonstrated significant difficulties in accurately simplifying the bitwise and arithmetic operations within each case, as shown in Listing 18:
Listing 18: Instruction Substitution DeepSeek R1 final deobfuscation attempt
1unsigned int secret_function(unsigned int n) {
2switch(n & 3) { // Original comparisons -> n%4
3case 0: // Block 0x27
4return ((0xBAAAD0BF | n) ^ (0xE6C98769 & ~n) ^ 2) *
5((n & ((0xBAAAD0BF | n) ^ (0xE6C98769 & ~n))) |
6(~n & 2));
7case 1: { // Block 0xb7
8unsigned int t = (~n | 0xBAAAD0BF) & 0x8ABD1CD5;
9return t * (n - 3);
10}
11case 2: { // Block 0x113
12unsigned int mask = 0x00F603E35; // Extended to 0xF603E35
13unsigned int a = (n & 0xBAAAD0BF) | (~n & mask);
14unsigned int b = (n ^ 0x811D5B51 | 0x7EE2A4AE) & 0x811D5B51;
15return a * b;
16}
17default: { // Block 0x1ac (case 3)
18unsigned int magic = 0xF4577A2A;
19unsigned int math = -(n + 0xBAAAD0BF); // Explicit wraparound
20unsigned int bits = (n | ~n ^ 5) & magic;
21return math * bits;
22}
23}
Comparing the output in Listing 18 with the reference implementation reveals substantial inaccuracies across all cases: Case 0: The reference implementation uses the simple expression (n | 0xBAAAD0BF) * (2 ^ n), but DeepSeek R1 produced an unnecessarily complex expression involving multiple nested operations: ((0xBAAAD0BF | n) ^ (0xE6C98769 & ~n) ^ 2) * ((n & ((0xBAAAD0BF | n) ^ (0xE6C98769 & ~n))) | (~n & 2)), introducing extraneous constants and operations not present in the original code. Case 1: While closer to the reference, DeepSeek R1 still unnecessarily complicated the expression with (~n | 0xBAAAD0BF) & 0x8ABD1CD5 instead of the simpler (n & 0xBAAAD0BF). The (n - 3) component is correctly identified but the overall expression remains needlessly complex. Case 2: DeepSeek R1 significantly diverged from the reference with a multi-step computation involving temporary variables and unrelated constants. The expression (n ^ 0x811D5B51 | 0x7EE2A4AE) & 0x811D5B51 is particularly problematic as it introduces constants not present in the original code and fails to capture the simple (n ^ 0xBAAAD0BF) * (4 | n) logic. Case 3: The model incorrectly represented the logic as -(n + 0xBAAAD0BF) * ((n | ~n ^ 5) & 0xF4577A2A) instead of the correct (n + 0xBAAAD0BF) * (5 & n), correctly identifying the addition operation but introducing sign inversion and unnecessarily complex bitwise operations. Interestingly, when presented with isolated bitwise expressions outside the context of assembly deobfuscation, DeepSeek R1 was able to simplify them correctly, suggesting the model’s difficulty stemmed from the contextual complexity of assembly analysis rather than an inherent inability to handle bitwise operations. DeepSeek R1 failed to successfully deobfuscate the assembly due to inaccurate simplification of bitwise and arithmetic operations despite identifying the switch structure.
3.2.5 Claude 3.7 Sonnet
Unlike other models, Claude 3.7 Sonnet took a distinctive approach by completely removing all the obfuscation constants and deriving a highly simplified mathematical pattern as shown in Listing 19:
Listing 19: Instruction Substitution Claude final deobfuscation attempt
1int function(int x) {
2int remainder = x & 3; // equivalent to x % 4
3int result;
4switch(remainder) {
5case 0:
6// The XOR/AND/OR sequence constructs (x ^ 2)
7// -1 ^ 2 = -3
8result = (-3) * (x ^ 2);
9break;
10case 1:
11// All the bit manipulations simplify to this
12result = x * (x + 3);
13break;
14case 2:
15// The XOR/AND/OR sequence constructs (x ^ 4)
16// -1 ^ 4 = -5
17result = (-5) * (x ^ 4);
18break;
19case 3:
20// Similar bit manipulations lead to this
21result = x * (x + 5);
22break;
23}
24return result;
25}
Claude’s approach in Listing 19 stands in stark contrast to other models, which preserved the original obfuscation constants (0xBAAAD0BF, 0xE6C98769, 0x8ABD1CD5, 0xF603E35, 0xF4577A2A) in their deobfuscated code. Claude instead identified an elegant mathematical pattern:
  • For even remainders (0, 2): result = (-1-remainder) * (x ^ (remainder+2))
  • For odd remainders (1, 3): result = x * (x + (remainder+2))
Claude’s approach in Listing 19 differs from other models by replacing the original obfuscation constants with a simple mathematical pattern based on remainder values. While this higher level of abstraction elegantly simplifies the code, it remains unverified whether this transformation preserves the exact behavior of the original assembly across all inputs without empirical validation.
3.3 Control flow flattening
In this section, we analyze the type known as Control flow flattening (CFF). This type of obfuscation aims at obfuscating the control flow by flattening the control flow graph. For a comprehensive description and additional details regarding this obfuscation technique, refer to [36, 44]. In this section we ignore several models like Grok 3, Chat o1, and ChatGPT 4.5 which successfully deobfuscated the code by correctly identifying the flattened switch structure and accurately reconstructing the computational logic without requiring additional guidance.
3.3.1 Grok 2
Grok 2 produced the following code as its final deobfuscation attempt Listing  20:
Listing 20: Control flow flattening Grok 2 final deobfuscation attempt
1int deobfuscated_function(int input) {
2 int flag = input & 3;
3 int result;
4 switch (flag) {
5 case 0: // flag == 0
6 result = (input + 0xbaaad0bf) * (input & 5);
7 break;
8 case 1: // flag == 1
9 result = (input | 0xbaaad0bf) * (input ^ 2);
10 break;
11 case 2: // flag == 2
12 result = (input ^ 0xbaaad0bf) * (input | 4);
13 break;
14 case 3: // flag == 3
15 result = (input & 0xbaaad0bf) * (input + 3);
16 break;
17 }
18 return result;
19}
Upon verification, Grok 2’s implementation in Listing 20 accurately identified the switch structure determined by the two least significant bits of the input (input & 3) and correctly derived each arithmetic and bitwise operation from the assembly. However, it incorrectly associated these derived expressions with their respective conditional paths, leading to mismatched logic in three out of four cases. In detail, the computation (input + 0xbaaad0bf) * (input & 5) correctly corresponds to the original logic for the condition input & 3 == 3, yet Grok 2 mistakenly assigned it to the condition input & 3 == 0. Similarly, the expression (input | 0xbaaad0bf) * (input ^ 2), originally intended for condition input & 3 == 0, was incorrectly assigned to condition input & 3 == 1. Additionally, Grok 2 placed the expression (input & 0xbaaad0bf) * (input + 3), correctly intended for input & 3 == 1, under condition input & 3 == 3. The only correct conditional assignment was the expression (input ^ 0xbaaad0bf) * (input | 4) correctly matched with the intended condition input & 3 == 2. The original obfuscation employed control flow flattening, using an intricate state machine initialized with the constant 0x64fd8a96. The assembly exhibited complex conditional jumps and redundant arithmetic instructions (e.g., repeated subtraction at addresses 0x46 and 0x5c), deliberately complicating control flow tracing. The presence of these complexities suggests challenges faced by Grok 2 in correctly mapping computational logic to conditional execution paths.
Grok 2 successfully isolated the correct arithmetic and bitwise transformations from heavily obfuscated code, demonstrating proficiency in arithmetic interpretation. Nonetheless, it showed a consistent pattern of errors in accurately associating these transformations with their corresponding control-flow conditions, specifically misaligning three out of four cases. Analysis reveals that Grok 2 failed to accurately deobfuscate the assembly code, primarily due to mismatched conditional logic, although it correctly identified the underlying arithmetic operations.
3.3.2 ChatGPT 4o
ChatGPT 4o made multiple attempts to deobfuscate the provided assembly. Its final deobfuscation attempt is presented in Listing 21:
Listing 21: Control flow flattening ChatGPT 4o final deobfuscation attempt
1int32_t obfuscated_function(uint32_t input) {
2 uint32_t mod_4 = input & 3;
3 uint32_t output;
4 switch (mod_4) {
5 case 0:
6 output = (input | 0xBAAAD0BF) * (input ^ 2);
7 break;
8 case 1:
9 output = (input & 0xBAAAD0BF) * (input + 3);
10 break;
11 case 2:
12 output = (input ^ 0xBAAAD0BF) * (input | 4);
13 break;
14 case 3:
15 output = (input + 0xBAAAD0BF) * (input & 5);
16 break;
17 }
18
19 return output;
20}
The final output from ChatGPT 4o shown in Listing 21 correctly identified the intended switch structure based on the input condition input & 3, accurately reproducing the arithmetic and bitwise transformations present in the original assembly logic. However, during its initial analysis, the model produced incorrect interpretations by misrepresenting the obfuscated assembly as a complex, nested conditional structure rather than recognizing the straightforward four-case logic dictated explicitly by input & 3. Initially, ChatGPT 4o introduced unnecessary intermediate state variables (such as 0xA4431FAA) and misidentified temporary computation values (e.g., 0xE9A524FA) as potential final outputs, instead of correctly tracing the final result stored at [ebp - 0x10].
After iterative prompting, explicitly clarifying that the control flow involved exactly four cases based on input & 3 and specifying the correct output variable ([ebp - 0x10]), ChatGPT 4o adjusted its interpretation and accurately represented the arithmetic and bitwise operations in their intended structure. ChatGPT 4o successfully deobfuscated the assembly with guidance but initially failed without it.
3.3.3 DeepSeek R1
The DeepSeek R1 initially interpreted the assembly as implementing a complex state machine with multiple transitions and conditional branches. In its first deobfuscation attempt, DeepSeek R1 produced a verbose pseudocode that, while technically accurate in tracing the state transitions, failed to identify the underlying simplicity of the algorithm. The model’s initial representation included unnecessary state variables and transition logic as shown in Listing 22:
Listing 22: Control flow flattening DeepSeek R1 initial attempt
1int obfuscated_function(int input) {
2int state = 0x64FD8A96;
3int result = 0;
4int mod = input & 3; // Equivalent to input % 4
5while(1) {
6 switch(state) {
7 case 0x64FD8A96:
8 state = (mod == 0) ? 0x707B18BC : 0xD5393A07;
9 break;
10 // Many more state transitions...
11 }
12}
13}
When prompted by the user with the simple instruction "knowing all that, simplify the method as much as possible," DeepSeek R1 immediately recognized the redundancy in its state machine representation and produced the fully simplified implementation shown in Listing 23:
Listing 23: Control flow flattening DeepSeek R1 final implementation
1int calculate(int input) {
2const int C = 0xBAAAD0BF; // Obfuscation constant
3int mod = input % 4;
4switch(mod) {
5 case 0: return (input | C) * (input ^ 2);
6 case 1: return (input & C) * (input + 3);
7 case 2: return (input ^ C) * (input | 4);
8 case 3: return (input + C) * (input & 5);
9}
10return 0; // Unreachable
11}
DeepSeek R1 showed a solid grasp of the assembly code’s core algorithm in Listing 23, correctly identifying all four cases based on input modulo 4 and their corresponding operations. Overall it successfully deobfuscated the assembly with prompting but initially failed to simplify it autonomously.
3.3.4 Claude 3.7 Sonnet
Claude demonstrated exceptional deobfuscation capabilities by correctly identifying the core algorithm on its first attempt. The model accurately recognized that the obfuscated assembly implemented a simple switch structure based on input modulo 4, successfully cutting through the complex state machine and control flow obfuscation, as shown in Listing 24:
Listing 24: Control flow flattening Claude final deobfuscation attempt
1unsigned obfuscated_calculation(unsigned param) {
2unsigned result = 0;
3const unsigned magic = 0xbaaad0bf; // Critical fix here
4switch(param % 4) {
5case 0:
6result = (param | magic) * (param ^ 2);
7break;
8case 1:
9result = (param & magic) * (param + 3);
10break;
11case 2:
12result = (param ^ magic) * (param | 4);
13break;
14case 3:
15result = (param + magic) * (param & 5);
16break;
17}
18return result;
19}
As shown in Listing 24, Claude 3.7 Sonnet successfully deobfuscated the assembly but introduced a minor constant error which was later fixed. Specifically, Claude initially misidentified the magic constant value (0xbaaad0bf) instead of (0xbaad0bf), which was subsequently corrected to ensure the deobfuscated code would produce identical results to the original assembly.
3.4 Combined Techniques
In this section, we analyze Combined Techniques (All), meaning bogus control flow, instruction substitution and control flow flattening applied to the same code. For a comprehensive description and additional details regarding this obfuscation technique, refer to [36, 44]. We are not testing weaker models like GPT-4o, GPT-3o Mini, and Grok2 here due to their particularly poor overall performance.
3.4.1 ChatGPT 4.5
In the initial analysis ChatGPT 4.5 misrepresented the assembly as a looping state machine involving a perpetual loop (while(1)) combined with a switch statement. However, manual verification confirmed that the actual assembly implements a single-pass switch structure beginning at address 0x156b, without looping behavior. Additionally, the assembly includes reads from address [0] at instructions 0x14 and 0x1b, consistently returning zero at runtime and significantly simplifying the logic. Despite this simplification, ChatGPT 4.5 initially treated redundant obfuscation instructions, such as jmp 0x7d8, as indicators of functional loops.
In a subsequent attempt, ChatGPT 4.5 correctly identified the high-level control structure based on the expression input & 3. However, the model erroneously assigned arbitrary constants (0xd87b0953, 0x3e2a6a0d), disregarding explicit computational sequences present after address 0x1546. Specifically, at address 0x1585, assembly instructions explicitly reference the constant 0xBAAAD0BF, yet ChatGPT 4.5 overlooked these operations by prematurely concluding its analysis based on intermediate jump table assignments observed at address 0x1579.
In its final attempt, ChatGPT 4.5 produced the C implementation shown in Listing 25:
Listing 25: All obfuscation final deobfuscation attempt 4.5
1unsigned int function(unsigned int input) {
2 unsigned int result;
3 switch (input & 3) {
4 case 0:
5 result = ((input & 0xBAAAD0BF) | (input ^ 0xBAAAD0BF)) * (~input & 2);
6 break;
7 case 1:
8 result = input - 1;
9 break;
10 case 2:
11 result = (~input ^ 0xBAAAD0BF) * (input + 0xAA61A945);
12 break;
13 case 3:
14 result = input - 2;
15 break;
16 }
17 return result;
18}
This final implementation in Listing 25 deviated substantially from the logic verified in the original assembly. For instance, the assembly at address 0x1585 clearly illustrates operations such as:
 mov esi, edx and esi, 0xbaaad0bf or esi, edx
This sequence corresponds explicitly to (input & 0xBAAAD0BF) | (input ^ 0xBAAAD0BF), yet ChatGPT 4.5 incorrectly introduced an unnecessary NOT operation ( input) and an inappropriate multiplier ( input & 2). Similarly, the assembly logic at case 1, representing the operation (input & 0xBAAAD0BF) * (input + 3), was oversimplified by ChatGPT 4.5 into merely input - 1, entirely omitting the required arithmetic and bitwise steps.
Furthermore, at 0x15e9, the verified assembly correctly implements (input ^ 0xBAAAD0BF) * (input | 4), but ChatGPT 4.5’s final interpretation introduced an extraneous constant (0xAA61A945) and a NOT operation, thus deviating significantly from the verified instructions. Lastly, for the computation explicitly performed at 0x1644(input + 0xBAAAD0BF) * (input & 5)—ChatGPT 4.5 simplified incorrectly to input - 2, entirely omitting essential arithmetic and bitwise operations explicitly evident in the assembly instructions.
Overall ChatGPT 4.5 failed to accurately deobfuscate the assembly despite recognizing the switch structure, producing incorrect arithmetic and bitwise operations.
3.4.2 ChatGPT-pro-o1
In its initial attempt, ChatGPT o1-proincorrectly represented the assembly logic as a looping state machine, as illustrated in Listing 26:
Listing 26: All obfuscation initial o1-pro attempt
1while (1) {
2 switch (local_input) {
3 case 0x8f77bef6: local_state = 0x8d270314; break;
4 case 0x8f77bef7: local_state = 0xc26f4570; break;
5 // Many more cases
6 }
7}
In this representation shown in Listing 26, o1-pro incorrectly inferred a persistent loop, misled by jump instructions (such as from 0x170d to 0x6b). The actual assembly, however, executes a single-pass switch statement based explicitly on input & 3 starting at address 0x156b. Additionally, reads from memory location [0] consistently yield zero, simplifying the analysis by resolving certain conditional checks at runtime. In subsequent iterations, o1-pro correctly identified the single-pass switch based on input & 3, yet incorrectly interpreted intermediate static constant assignments as final computational results, illustrated in Listing 27:
Listing 27: All obfuscation intermediate result o1-pro
1switch (input & 3) {
2 case 0: result = 0xd87b0953; break;
3 case 1: result = 0x35a94b5b; break;
4 case 2: result = 0xecdd7d46; break;
5 case 3: result = 0xfd7cd41e; break;
6}
These static assignments in Listing 27 (e.g., mov dword ptr [ebp - 0x50], 0xd87b0953 at 0x1579) are components of obfuscation noise rather than definitive outputs. Critical computational logic, such as arithmetic operations involving constants like 0xBAAAD0BF, appear later in the assembly beyond address 0x1546, specifically in code blocks at addresses 0x1585, 0x15e9, 0x165b, and 0x16a7. ChatGPT o1-pro consistently neglected these dynamic computational sequences. The final deobfuscation attempt by o1-pro produced the incorrect approximation shown in Listing 28:
Listing 28: All obfuscation final deobfuscation attempt o1-pro
1unsigned int function(unsigned int input) {
2 unsigned int transformed = 0;
3 switch (input & 3) {
4 case 0: transformed = ((input ^ 0xBAAAD0BF) & 0xBAAAD0BF) | ((input & ~(input ^ 2)) ^ 2); break;
5 case 1: transformed = ((~input ^ 0xBAAAD0BF) & (~input | (0 - 3))) + (~input & (0 - 3)); break;
6 case 2: transformed = (((input ^ 0xBAAAD0BF) & ~(input ^ 0xBAAAD0BF)) ^ ~(input ^ 4)) & 4; break;
7 case 3: transformed = (0 - 0xBAAAD0BF - input) & (input & 5); break;
8 }
9 return transformed - 1;
10}
Although this version in Listing 28 recognized the single-pass switch, the arithmetic and bitwise transformations differed significantly from the assembly’s actual logic. At address 0x1585, the assembly explicitly computes (input & 0xBAAAD0BF) | (input ^ 0xBAAAD0BF) followed by a multiplication involving input & 2, yet o1-pro introduced unrelated bitwise combinations without the necessary arithmetic multiplication. Similar inaccuracies occurred at addresses 0x15e9, 0x165b, and 0x16a7, where essential arithmetic multiplications and logical operations involving constants like 0xBAAAD0BF were incorrectly replaced by o1-pro with oversimplified bitwise logic.
Overall ChatGPT-pro-o1 failed to accurately deobfuscate the assembly despite identifying the switch structure, misrepresenting key arithmetic operations.
3.4.3 Grok 3
Initially, Grok 3 incorrectly modeled the assembly code as a looping state machine with constant states, exemplified by the pseudocode in Listing 29:
Listing 29: All obfuscation initial step Grok 3
1function process_input(input):
2 state = 0x4c3945a8
3 while true:
4 if state == 0x4c3945a8:
5 if input < 10:
6 state = 0xd87b0953
7 else if state == 0xd87b0953:
8 return 0xd87b0953
In reality, the assembly implemented a single-pass switch structure at address 0x156b, determined solely by input & 3. The model’s misunderstanding stemmed primarily from misleading jump instructions and constant-value comparisons (e.g., sub eax, 0x8f77bef6), as well as consistent zero-value memory reads from address [0]. Subsequently, Grok 3 further simplified its interpretation to a constant return function:
int simplified_function() {
    return 0xd87b0953;
This version entirely neglected dynamic computations dictated by input & 3, incorrectly assuming a static result. In a later attempt, Grok 3 correctly identified the underlying switch-based logic at address 0x156b but inaccurately represented computational operations with generic placeholders, as demonstrated in Listing 30:
Listing 30: All obfuscation intermediate step Grok 3
1unsigned int function(unsigned int input) {
2 unsigned int result;
3 switch (input & 3) {
4 case 0: result = (input ^ 0x12345678) + 0x9ABCDEF0; break;
5 case 1: result = (input & 0xFEDCBA98) | 0x13579BDF; break;
6 case 2: result = (input << 2) | (input >> 30); break;
7 case 3: result = input * 0x2468ACE0; break;
8 }
9 return result;
10}
This code in Listing 30 significantly deviated from the actual arithmetic operations present in the assembly, particularly those involving constants like 0xBAAAD0BF and explicit arithmetic multiplications (imul). In its final iteration, Grok 3 presented the implementation shown in Listing 31:
Listing 31: All obfuscation final step Grok 3
1unsigned int function(unsigned int input) {
2 unsigned int result;
3 switch (input & 3) {
4 case 0: result = ((input & 0xBAAAD0BF) | (input ^ 0xBAAAD0BF)) * (~input & 2); break;
5 case 1: result = input - 1; break;
6 case 2: result = (~input ^ 0xBAAAD0BF) * (input + 0xAA61A945); break;
7 case 3: result = input - 2; break;
8 }
9 return result;
10}
Although the final version in Listing 31 correctly reflected the conditional structure based on input & 3, it deviated substantially from the precise computations in the original assembly. For case 0, Grok 3 introduced a bitwise negation ( input), an incorrect bitwise OR operation, and an improper multiplier, instead of accurately representing the original assembly logic: (input | 0xBAAAD0BF) * (input ^ 2). Similarly, in case 1, the actual computation (input & 0xBAAAD0BF) * (input + 3) was erroneously simplified to input - 1, omitting essential arithmetic operations. Case 2 further demonstrated inaccuracies, as Grok 3 incorrectly utilized a bitwise negation and introduced an unrelated constant 0xAA61A945, diverging from the intended computation (input ^ 0xBAAAD0BF) * (input | 4). Finally, case 3 was significantly oversimplified to input - 2, failing to include the necessary arithmetic addition and bitwise AND: (input + 0xBAAAD0BF) * (input & 5). Overall Grok 3 failed to accurately deobfuscate the assembly despite identifying the switch structure, misrepresenting key arithmetic and logical operations.
3.4.4 DeepSeek R1
DeepSeek R1’s deobfuscation attempts on the provided assembly code revealed significant limitations in handling complex obfuscation techniques. The initial misconception is similar to other models analyzed, DeepSeek R1 incorrectly interpreted the assembly as implementing a state machine with continuous execution as shown in Listing 32:
Listing 32: DeepSeek R1 initial interpretation
1# State machine with value-controlled transitions
2state = 0x4C3945A8 # Initial state
3while True:
4 # State transitions based on complex condition checks
5 if state == 0x4C3945A8:
6 if (condition1 and condition2):
7 state = 0xD87B0953
8 else:
9 state = 0x8D270314
10 # ...other state transitions...
This interpretation in Listing 32 fundamentally misunderstood the code’s structure, which actually implemented a single-pass switch statement determined by input & 3 at address 0x156b, not a continuous-execution state machine.
Despite additional guidance, DeepSeek R1 failed to accurately identify the critical computation blocks that follow address 0x1546. When directed to focus on the switch structure based on edi & 3, the model produced the code shown in Listing 33:
Listing 33: DeepSeek R1 switch implementation
1uint32_t calculate_result(uint32_t input) {
2 // State machine with 4 primary execution paths
3 uint32_t state = 0x4C3945A8;
4 uint32_t a = 0, b = 0;
5
6 // Core switch structure
7 uint32_t temp1, temp2;
8
9 switch(input & 3) {
10 case 0:
11 temp1 = (input ^ 0xBAADF00D) << 3;
12 temp2 = (input + 0x715BBD7F) | 0xDEADBEEF;
13 break;
14 // ...other cases...
15 }
16
17 // Common final processing
18 uint32_t result = temp1 * temp2;
19 result ^= state;
20 result = (result >> 16) | (result << 16);
21 return result ^ 0x35DCA7D6;
22}
Perhaps most concerning was DeepSeek R1’s introduction of fictional hexadecimal values nowhere present in the original assembly like: 0xBAADF00D, 0xDEADBEEF, 0xCAFEBABE and others, as seen in Listing 33. These are recognizable “hexspeak” values (hexadecimal values that resemble English words) commonly used as placeholders in debugging or illustrative examples, but were entirely fabricated by the model rather than derived from the assembly code. When challenged about the origin of these values, DeepSeek R1 acknowledged they were not present in the actual assembly but provided no meaningful correction to its analysis. DeepSeek R1 failed to deobfuscate the assembly, misinterpreting the structure and fabricating constants not present in the code.
3.4.5 Claude 3.7 Sonnet
Claude attempted to deobfuscate the provided assembly code across multiple iterations. The model’s approach evolved through several distinct phases, each demonstrating specific limitations in handling complex obfuscated code. In its initial analysis, Claude characterized the assembly as implementing a state machine with the structure shown in Listing 34:
Listing 34: Claude initial state machine analysis
1int function(int input) {
2 // Initial setup
3 bool isEven = (input % 2 == 0);
4 bool isLessThan10 = (input < 10);
5 int lastTwoBits = input & 3;
6 int result = 0;
7
8 // Different computation paths based on input properties
9 if (isEven && isLessThan10) {
10 result = compute_value_1(input, lastTwoBits);
11 } else if (!isEven && isLessThan10) {
12 result = compute_value_2(input, lastTwoBits);
13 } else if (isEven && !isLessThan10) {
14 result = compute_value_3(input, lastTwoBits);
15 } else {
16 result = compute_value_4(input, lastTwoBits);
17 }
18
19 return result;
20}
When prompted to elaborate on each computation path, Claude expanded its analysis but represented these operations as simple lookup tables rather than preserving the arithmetic and bitwise logic present in the assembly. This intermediary analysis revealed a pattern-matching approach focused on detecting branching conditions rather than precise computation. When specifically directed to use a switch structure based on input & 3, Claude produced its final deobfuscation attempt shown in Listing 35:
Listing 35: Claude final deobfuscation attempt
1int function(int input) {
2bool isLessThan10 = (input < 10);
3int lastTwoBits = input & 3;
4switch (lastTwoBits) {
5 case 0: // Even number, divisible by 4
6 return isLessThan10 ? 4 : 0;
7 case 1: // Odd number, one more than multiple of 4
8 return isLessThan10 ? 3 : 2;
9 case 2: // Even number, 2 more than multiple of 4
10 return isLessThan10 ? 4 : 1;
11 case 3: // Odd number, 3 more than multiple of 4
12 return isLessThan10 ? 5 : 3;
13}
14
15return -1; // This should never happen
16}
This implementation in Listing 35 fails to capture critical aspects of the original assembly: the constant 0xBAAAD0BF that appears at addresses 0x598, 0xa18, and elsewhere is entirely absent from Claude’s representation, despite its central role in the bitwise operations. The original code at address 0x1585 contains operations like (input & 0xBAAAD0BF) | (input ^ 0xBAAAD0BF), which are completely absent from Claude’s deobfuscation. The assembly clearly shows multiplication operations (e.g., at 0x1392-0x13b1) that form essential components of the computation, yet these are not represented in Claude’s output. Claude represented the function as returning fixed integer values (0-5) based solely on the input’s properties, rather than capturing the dynamic computations performed by the assembly code. Additionally, Claude’s representation eliminated all intermediate calculations and variable manipulations present in the original assembly, significantly altering the function’s actual behavior. Claude 3.7 Sonnet failed to accurately deobfuscate the assembly, missing key arithmetic and bitwise operations despite identifying the switch structure.
4 Discussion
Our systematic evaluation of several state-of-the-art LLMs across multiple obfuscation techniques reveals significant performance variations, from autonomous deobfuscation to complete failure, depending on the technique employed. Rather than attributing these differences to singular factors like model size or architecture, we propose a theoretical framework based on four distinct yet interconnected dimensions: Reasoning Depth (analytical processing of program logic), Pattern Recognition (structural identification amid obfuscation), Noise Filtering (distinguishing essential logic from deceptive constructs), and Context Integration (maintaining coherence across fragmented code). This framework provides definitive explanatory power for our results and establishes predictive value for future advances. The following subsections explore each dimension through established program analysis theory, analyze cybersecurity implications, address study limitations, and identify promising research directions for enhancing AI-assisted reverse engineering.
4.1 Theoretical Framework for LLM Deobfuscation
The application of large language models to assembly-level deobfuscation represents a novel intersection of natural language processing and program analysis, requiring a structured approach to understand their capabilities and limitations. Our four-dimensional framework, derived from pattern analysis across multiple obfuscation techniques and models,provides a systematic foundation for characterizing how LLMs process obfuscated code and why specific techniques present greater analytical challenges than others.
Context Integration Dimension
Context Integration refers to a model’s ability to maintain coherence across logically related but physically disconnected code segments. This dimension parallels control flow graph analysis [54] and leverages transformer architecture’s self-attention mechanisms [4], which capture long-range dependencies across token sequences. Our experiments revealed varied capabilities: GPT-4.5, GPT-o1, and Grok 3 demonstrated superior integration, successfully reconstructing relationships between computations at address 0x1585 with condition input & 3 == 0 (Section 3.3), despite fragmentation through state machines initialized with constant 0x64fd8a96. Claude 3.7 Sonnet correctly determined that jumps to address 0x40169e were unreachable due to opaque predicates but struggled with control flow flattening (Section 3.1). DeepSeek R1 and GPT-3o Mini produced state machine representations with transitions like state = 0x4c3945a8 to state = 0xd87b0953 (Section 3.4.4 and Section 3.1.4) without simplifying to underlying algorithms. When facing combined obfuscation, all models struggled, with even GPT-4.5 misinterpreting intermediate state values at 0x1579 as definitive outputs (Section 3.4.1). We identified a clear relationship between context window size and performance, models with windows exceeding 200,000 tokens demonstrated superior performance on complex tasks.
Reasoning Depth Dimension
Reasoning Depth reflects a model’s capacity to perform logical inference on program properties similar to formal verification and abstract interpretation [46], but through data-driven approaches rather than explicit logical engines. This capability is based on chain-of-thought reasoning [39], allowing models to decompose complex problems into manageable steps. Our analysis revealed a clear performance spectrum: Claude 3.7 Sonnet and GPT-4.5 demonstrated advanced reasoning, correctly determining that expressions like ((var1 - 1) * var1) & 1 == 0 always evaluate to true (Section 3.1) — effectively recognizing mathematical invariants without guidance. GPT-Pro-o1 and Grok 3 showed moderate capabilities (Section 3.1.3 and Section 3.1.5), requiring occasional hints, while GPT-4o, DeepSeek R1, and GPT-3o Mini exhibited limited reasoning (Section 3.1.1, Section 3.1.6, and Section 3.1.4), often treating invariant conditions as variable and missing unreachable code paths. This gradient underscores reasoning depth as a critical capability bridging general language processing and domain-specific program analysis.
Pattern Recognition Dimension
Pattern Recognition involves identifying structural and computational patterns within obfuscated code, building upon program similarity metrics and clone detection theory [50, 51]. This dimension connects to the "naturalness of software" hypothesis [52], which proposed that programming languages contain statistical regularities enabling probabilistic modeling. Our experiments showed that models like GPT-4.5, GPT-o1, and Grok 3 successfully reconstructed the underlying switch structures despite extensive obfuscation (Section 3.3), identifying complex arithmetic transformations such as (input | 0xbaaad0bf) * (input ^ 2) within heavily obfuscated contexts (Section 3.2). This capability emerged organically from general language modeling rather than specialized program analysis tools. Performance varied significantly between obfuscation techniques - models that excelled at reconstructing control structures often struggled with arithmetic transformations, supporting the findings that statistical models of code can capture both syntactic and semantic patterns [53], with varying degrees of success.
Noise Filtering Dimension
Noise Filtering encompasses a model’s ability to distinguish essential computational logic from obfuscation artifacts, sharing functional similarities with program slicing [48] despite fundamentally different mechanisms. From an information-theoretic perspective [47], LLMs must extract meaningful signal from the increased entropy that obfuscation introduces. Our experiments demonstrated this capability in specific contexts: when analyzing bogus control flow, models like GPT-4.5, GPT-o1, and Claude 3.7 Sonnet correctly identified repetitive predicate checks as non-functional (Section 3.1.2, Section 3.1.3, Section 3.1), simplifying the code to its essential form. However, we observed significant limitations: DeepSeek R1 generated convoluted expressions like ((0xBAAAD0BF | n) ^ (0xE6C98769 & ~n) ^ 2) instead of the simpler (n | 0xBAAAD0BF) * (n ^ 2) (Section 3.2.4). This parallels challenges in adversarial machine learning [49], where obfuscation disrupts the statistical patterns on which LLMs rely, creating noise that confounds even models with strong capabilities in other dimensions.
4.2 Mapping Framework Dimensions to Empirical Results
Our four-dimensional framework provides explanatory power for the performance variations observed in Table 2. Each obfuscation technique challenges specific dimensional capabilities, creating predictable performance patterns across models.
Bogus Control Flow and Reasoning Depth
Performance against bogus control flow (BCF) appears to be associated with reasoning capabilities. Models with lower intervention requirements (Level 0-2) demonstrated superior reasoning depth by correctly analyzing opaque predicates like ((var1 * (var1 - 1)) & 1 == 0(Section 3.1). In contrast, models requiring expert intervention (Levels 4-5) consistently failed to recognize mathematical invariants, treating potentially-always-true conditions as variable. This dimension clearly separates models into performance tiers: high performers (Claude 3.7 Sonnet, GPT-4.5, Grok3), moderate performers (GPT-Pro-o1), and low performers (GPT-4o, GPT-3o Mini, DeepSeekR1).
Instruction Substitution and Pattern Recognition
Results for instruction substitution (IS) reveal widespread deficiencies in pattern recognition (Section 3.2). The consistently high intervention requirements demonstrate how obfuscated arithmetic operations challenge even advanced models’ pattern recognition capabilities. We identified various approaches: some models attempted to preserve obfuscation constants (Section 3.2.2, Section 3.2.4) while others sought mathematical simplification (Section 3.2.5). Neither approach consistently succeeded, with most models requiring significant intervention (Levels 2-5), highlighting pattern recognition as a critical bottleneck in current deobfuscation capabilities.
Control Flow Flattening and Context Integration
Performance against control flow flattening (CFF) appears to reflect context integration capabilities. Several models (GPT-4.5, GPT-Pro-o1, Grok3) achieved level 0 (Section 3.3), successfully maintaining coherence across fragmented code blocks and reconstructing relationships between computations at address 0x1585 and control conditions at 0x156b (Section 3.3.2). Most models performed relatively well on this task (Levels 0-2), suggesting that context integration is a comparative strength in current LLM architectures, potentially linked to attention mechanisms optimized for capturing long-range dependencies.
Combined Techniques and Dimensional Interdependence
The universal failure against combined techniques (All: Level 5 across all models tested in Section 3.4) demonstrates how obfuscation specifically targets the interdependence of these four dimensions. When multiple techniques simultaneously challenge reasoning depth, pattern recognition, noise filtering, and context integration, even models that perform well in individual dimensions falter. This suggests that dimensional capabilities operate synergistically rather than independently: A model strong in three dimensions but weak in one may still completely fail when all dimensions are challenged simultaneously.
Dimensional Performance Anomalies
Our analysis revealed important dimensional performance patterns that open new valuable research directions. For instance, Claude 3.7 Sonnet demonstrated exceptional Reasoning Depth with bogus control flow (Level 0) (Section 3.1) but showed comparatively weaker Context Integration with control flow flattening (Level 0-1) (Section 3.3.4), despite both techniques requiring sophisticated reasoning capabilities. Similarly, models with large context windows occasionally performed worse than smaller-context models on tasks that seemingly should benefit from extended context. These patterns demonstrate significant interactions between dimensional capabilities and specific implementation details of obfuscation techniques that our current framework does not fully capture. Future research should specifically investigate these boundary cases to refine our understanding of how dimensional capabilities interact and potentially identify additional factors that influence deobfuscation performance.
Dimensional Prioritization by Obfuscation Type
Our results suggest that each obfuscation technique challenges a different primary dimension:
  • Bogus Control Flow primarily challenges Reasoning Depth
  • Instruction Substitution primarily challenges Pattern Recognition
  • Control Flow Flattening primarily challenges Context Integration
  • Combined Techniques challenge all dimensions simultaneously
This dimensional mapping explains the inconsistent performance patterns observed across techniques: models excel where their dimensional strengths align with technique demands, while struggling where techniques target their dimensional weaknesses. Future architectural improvements should focus on balanced enhancement across all four dimensions rather than optimizing for any single capability.
4.2.1 Taxonomy of LLM Deobfuscation Errors
Analysis of deobfuscation attempts across our evaluation reveals distinct error patterns that occurred consistently across multiple models. The following taxonomy classifies these errors, with cross-references to examples already detailed in our Results section:
Predicate Misinterpretation Errors
These errors involve failing to recognize that certain conditions are always true or false (invariant), relating directly to limitations in the Reasoning Depth dimension of our framework. When analyzing bogus control flow, GPT-4o misinterpreted the opaque predicate (var1 * (var1 - 1)) & 1 == 0 as potentially false (Section 3.1.1), despite the mathematical property that the product of consecutive integers is always even. This misinterpretation led to incorrect control flow analysis, with the model treating the unreachable jump to address 0x40169e as a viable execution path. Similarly to the challenges faced in formal verification [46], these errors demonstrate the difficulty in symbolically reasoning about invariant properties. This error was prevalent in the analysis of bogus control flow obfuscation, appearing in 4 of 7 models (GPT-4o, GPT-3o Mini, Grok 2, DeepSeek R1 in Section 3.1). In contrast, models with enhanced reasoning capabilities, such as Claude 3.7 Sonnet, GPT-4.5, and Grok 3 (in Section 3.1), correctly identified the predicate’s invariant property. Beyond its frequency, this error has practical consequences, such as misidentification of potentially exploitable code paths in vulnerability analysis.
Structural Mapping Errors
These errors occur when a model correctly identifies computational components but incorrectly maps them to control structures, reflecting limitations in both the Pattern Recognition and Context Integration dimensions of our framework. When analyzing CFF - Grok 2 demonstrated this error type, correctly identifying four arithmetic transformations but assigning three to incorrect conditional paths. Specifically, it matched (input + 0xbaaad0bf) * (input & 5) to input & 3 == 0 instead of input & 3 == 3 (Section 3.3.1). This creates deobfuscated code that preserves computational elements but misimplements their control dependencies, meaning that the conditions governing their execution are wrong, resulting in runtime behavior fundamentally different from the original despite structural similarity. This type of error presents a particularly challenging issue for code analysis applications, where incorrect mappings may pass superficial validation, yet introduce subtle semantic errors that fundamentally alter program behavior[53]. We observed this error in 3 of 7 models (Grok 2, DeepSeek R1, GPT-4o in Section 3.3), while models with stronger context integration capabilities, such as larger context windows or advanced attention mechanisms (e.g., GPT-4.5 and Claude 3.7 Sonnet in Section 3.3), correctly reconstructed these relationships, particularly when dealing with less complex obfuscation techniques.
Control Flow Structure Misinterpretation Errors
This error type, primarily reflecting limitations in our framework’s Noise Filtering and Pattern Recognition dimensions, occurs when models incorrectly reconstruct the fundamental control flow structure of the original code. A common manifestation we observed was the incorrect inference of iterative structures (loops, state machines) where only single-pass logic exists. When analyzing bogus control flow obfuscation, DeepSeek R1 and GPT-3o Mini misinterpreted repetitive predicate checks (deliberately inserted as obfuscation artifacts) as indicators of functional looping behavior. For example, DeepSeek R1 generated for_n range(10) constructs in its deobfuscated output, despite the original assembly implementing a non-iterative execution path (Section 3.1.6). Similarly, when facing combined obfuscation techniques, ChatGPT-pro-o1 and Grok 3 reconstructed the code as a looping state machine (e.g., while (1) switch (state) ... (Section 3.4.2 and Section 3.4.3)) rather than identifying the single-pass switch statement at address 0x156b. While this error resembles hallucination phenomena in other LLM applications [55], our evidence demonstrates it primarily arises from misclassifying obfuscation patterns. These patterns, such as control flow flattening, transform direct branches into state-machine-like constructs to mislead static analysis tools. This misinterpretation carries significant security implications, as it can fundamentally alter an analyst’s understanding of program execution, potentially obscuring vulnerabilities or introducing nonexistent execution paths. These consequences directly impact vulnerability assessment procedures, where accurate control flow reconstruction is essential to identify potentially exploitable paths [56, 57], particularly in contexts where automated tools supplement human analysis. We observed this error in 5 out of 7 models (DeepSeek R1, GPT-3o Mini, ChatGPT-pro-o1, Grok 3, Claude 3.7 Sonnet), highlighting a widespread difficulty in separating obfuscation from actual control flow.
Arithmetic Transformation Errors
These errors involve failing to correctly reconstruct precise arithmetic and bitwise operations from obfuscated assembly, directly reflecting limitations in the Pattern Recognition dimension of our framework. All seven models exhibited this error type when confronting instruction substitution techniques, though with varying severity. Our analysis documented systematic pattern failures: ChatGPT-4o (Section 3.2.1) replaced the operation (input | 0xBAAAD0BF) * (input ^ 2) with entirely incorrect expressions containing unrelated constants and inappropriate operations. GPT-4.5 demonstrated a different failure mode (Section 3.2.2), preserving core operations but introducing unnecessary complexity by adding extraneous bitwise operations and constants to what should be simple expressions. For example, instead of the straightforward operation present in the original code, it generated (( input ^ 0xBAAAD0BF ^ 0xE6C98769) + 2) * (input | 2) . DeepSeek R1 (Section 3.2.4) produced particularly convoluted transformations with multiple nested operations and unrelated constants. Models adopted different strategies when handling obfuscated arithmetic operations. Claude 3.7 Sonnet (Section 3.2.5) attempted mathematical abstraction, deriving patterns like (-3) * (x ^ 2) for even-indexed cases while ignoring specific constants. GPT-4.5 preserved core operations but introduced unnecessary complexity. Grok 3(Section 3.2.3) achieved partial accuracy, correctly reconstructing some operations while misrepresenting others. DeepSeek R1 produced the most convoluted transformations with multiple nested operations and fabricated constants. These diverse approaches, each with distinct trade-offs between abstraction and fidelity, demonstrate the fundamental challenge LLMs face when attempting to simplify deliberately obfuscated arithmetic. The consistency of these errors across all models when faced with instruction substitution, a technique specifically designed to obscure individual operations, suggests a fundamental limitation in the ability of current LLMs to recognize equivalent computational expressions, particularly when obfuscation deliberately transforms simple operations into more complex forms through mathematical identities (all examples from Section 3.2).
Constant Propagation Errors
These errors involve incorrectly handling, identifying, or fabricating literal values in deobfuscated code, revealing limitations in both Pattern Recognition and Context Integration dimensions. We identified three distinct manifestations with varying severity. First, subtle transcription errors occurred when models processed large hexadecimal constants, exemplified by Claude 3.7 Sonnet’s (Section 3.2.5) misreading of 0xBAAAD0BF as 0xBAAAAD0BF (inserting an extra ’A’) during instruction substitution analysis. Second, models introduced mathematically unrelated constants, as seen when ChatGPT-4.5 (Section 3.2.2) generated expressions containing values like 0xE6C98769 and 0x8ABD1CD5 — constants not present in the original assembly but erroneously incorporated into arithmetic operations. Most concerning was the third category: complete fabrication of well-known "hexspeak" values. When analyzing combined obfuscation techniques, DeepSeek R1 generated deobfuscated code containing entirely fictional constants (Section 3.4.4) like 0xBAADF00D, 0xDEADBEEF, and 0xCAFEBABE—recognizable placeholder values commonly used by developers but completely absent from the original assembly. This fabrication suggests a problematic pattern-matching behavior where models revert to familiar training examples rather than faithfully representing the analyzed code. From a security perspective, these errors are particularly troubling when analyzing cryptographic implementations or any code where precise constant values directly impact functionality and correctness, such as protocol implementations, hash functions, or file format parsers. We identified constant propagation errors in 3 of 7 models (Claude 3.7 Sonnet, DeepSeek R1, GPT-4.5 in (Section 3.2)), indicating that even advanced LLMs struggle with accurate representation of symbolic information.
4.2.2 Obfuscation Technique Resistance Model
Based on our empirical findings, we propose a three-tier resistance model that categorizes obfuscation techniques according to their effectiveness against LLM-based deobfuscation:
Low Resistance Techniques
Bogus control flow obfuscation primarily challenges the Reasoning Depth dimension through opaque predicates and unreachable code paths. Our experiments revealed clear performance stratification: top performers (Claude 3.7 Sonnet, GPT-4.5, Grok 3) required minimal or no guidance (Levels 0-2), while others needed expert intervention or failed entirely. This stratification directly correlates with each model’s ability to perform mathematical reasoning about invariant conditions. Models with stronger reasoning capabilities performed consistently better (Section 3), demonstrating that this obfuscation technique, traditionally effective against static analysis tools, offers limited protection against advanced LLMs with enhanced reasoning capabilities.
Moderate Resistance Techniques
Control flow flattening primarily challenges the Context Integration dimension by fragmenting logically connected code segments across a state machine structure. This technique demonstrated intermediate resistance, with only two models achieving Level 0-1 success. Most models either required substantial guidance or failed entirely, frequently producing complex state machine representations that failed to simplify to the underlying algorithm. Effective deobfuscation requires both strong pattern recognition to identify computational elements and robust context integration to correctly reconstruct their relationships—a combination present in fewer models than those capable of handling bogus control flow.
High Resistance Techniques
Instruction substitution and combined obfuscation techniques demonstrated the highest resistance, with all models either requiring expert intervention (Level 5) or failing completely. Instruction substitution specifically challenges Pattern Recognition by replacing simple operations with complex mathematical equivalents. Combined techniques simultaneously attack multiple capability dimensions, creating a compounding effect that overwhelmed even the most capable models. Every model exhibited multiple error types from our taxonomy when facing combined obfuscation(Section 3.4). The universal failure against these techniques demonstrates a current upper bound on LLM deobfuscation capabilities, regardless of model size or architecture.
4.3 Limitations
Our systematic approach incorporates deliberate scoping decisions that define clear boundaries for this research:
Sample Diversity
Our evaluation strategically focused on a well-documented program obfuscated with OLLVM, limiting generalizability. While this controlled approach enabled detailed analysis across models, real-world malware employs diverse obfuscation tools and techniques. Future work should expand testing across multiple codebases and obfuscation frameworks to validate whether our dimensional framework applies consistently across varied obfuscation implementations.
Model Selection
Our study focused on seven commercial LLMs representing current state-of-the-art capabilities but cannot capture the full landscape. Notably absent are domain-specific code models and open-source alternatives that might exhibit different capabilities. This limitation reflects the practical constraints of comprehensive evaluation rather than intentional exclusion. Future research should evaluate specialized code-oriented models that may demonstrate different dimensional strengths.
Qualitative Assessment
Our attacker knowledge level framework, while providing structured evaluation, inherently involves subjective judgment. To mitigate this limitation, we conducted multiple deobfuscation attempts per scenario and documented detailed reasoning for each assigned level. However, the framework’s reproducibility remains a challenge that future work should address through developing more objective, quantitative metrics for deobfuscation success.
Temporal Boundaries
This work establishes a methodological foundation that will remain relevant despite the rapid evolution of LLM capabilities, with significant advances occurring on timescales of months. Our evaluation captures capabilities at a specific point (March 2025), establishing a baseline against which future progress can be measured. By focusing on dimensional capabilities rather than specific implementations, our framework provides enduring value for evaluating emerging models and techniques.
4.4 Future Research Directions
Our dimensional framework and empirical findings reveal several high-impact research opportunities that could advance both LLM-based deobfuscation capabilities and security defenses:
Dimensional Transfer Learning
Our results demonstrate uneven development across capability dimensions, with models exhibiting strength in specific dimensions (e.g., Context Integration) while struggling in others (e.g., Pattern Recognition). Future research should investigate whether techniques that enhance performance in one dimension could transfer to others. For example, research could explore whether pre-training approaches that improve reasoning capabilities for mathematical invariants (addressing Predicate Misinterpretation Errors) might also enhance pattern recognition for arithmetic transformations. Such dimensional transfer could provide a more efficient pathway to developing balanced capabilities across all four dimensions than treating each as an isolated challenge.
Adversarial Obfuscation Framework
The error taxonomy we’ve documented provides a foundation for developing next-generation obfuscation techniques specifically designed to resist LLM-based analysis. Future research should systematically explore techniques that exploit identified weaknesses across multiple dimensions simultaneously. Particularly promising are techniques that:
  • Embed mathematically complex opaque predicates that appear variable but are invariant (targeting Reasoning Depth)
  • Employ instruction substitution with mathematical identities that preserve human readability while maximizing LLM confusion (targeting Pattern Recognition)
  • Introduce structurally ambiguous constructs that suggest iterative execution but implement single-pass logic (targeting Noise Filtering)
  • Fragment related code blocks across non-obvious execution paths (targeting Context Integration)
These techniques could be implemented as extensions to existing obfuscation frameworks like OLLVM and systematically evaluated against both current and future LLM architectures.
Uncertainty-Aware Deobfuscation Systems
Our finding that all models struggled with precise arithmetic reconstruction suggests the need for systems that explicitly quantify uncertainty in their deobfuscation outputs. Future work should develop approaches where models provide confidence scores for different aspects of deobfuscated code, highlighting areas that may require human verification. This would transform deobfuscation from a binary success/failure paradigm to a collaborative human-AI workflow where resources are efficiently allocated to the most uncertain aspects of analysis. Such systems could potentially incorporate multiple specialized models, each focused on different dimensional capabilities, with a meta-model integrating their outputs while tracking uncertainty across the deobfuscation process.
Cross-Architecture Generalization
Our evaluation focused on x86_64 assembly, but real-world reverse engineering requires analyzing diverse architectures (ARM/AArch64, x86-32 [58], and emerging ISAs like, for example, RISC-V [59]). Future research should investigate whether the dimensional capabilities we have identified generalize across instruction set architectures, particularly exploring whether models trained primarily on one architecture can transfer capabilities to others. This research direction is particularly important given the increasing diversity of embedded systems and IoT devices that may employ obfuscation techniques. Methodologically, this would require developing standardized datasets of equivalent programs compiled and obfuscated across multiple architectures, enabling a direct comparison of deobfuscation performance.
Quantitative Benchmark Development
To enable rigorous tracking of progress in this field, future work should develop standardized quantitative benchmarks that objectively measure each dimension of our framework. Such benchmarks could include:
  • Reasoning depth metrics based on correct evaluation of increasingly complex opaque predicates
  • Pattern recognition metrics measuring accuracy in reconstructing obfuscated arithmetic operations
  • Noise filtering metrics quantifying ability to distinguish functional from non-functional code
  • Context integration metrics assessing accuracy in reconstructing control flow relationships
These benchmarks should include both synthetic examples designed to isolate specific capabilities and real-world obfuscated code samples to ensure practical relevance. By establishing such metrics, the field can move beyond qualitative assessments to measurable progress in LLM-based deobfuscation capabilities.
5 Conclusion
Our comprehensive evaluation of state-of-the-art LLMs on assembly-level deobfuscation establishes a new paradigm for understanding AI capabilities in cybersecurity contexts. This research delivers a robust theoretical framework, decisive evidence, and far-reaching implications that transform our understanding of automated code analysis.
The results demonstrate clear performance patterns across the obfuscation techniques. Several models deobfuscated bogus control flow with minimal intervention, while all models - regardless of architecture or size - failed against combined obfuscation techniques. Control flow flattening and instruction substitution presented intermediate challenges, with performance varying significantly across models. This performance gradient definitively maps the boundaries of current LLM capabilities for advanced binary analysis tasks.
The four-dimensional framework we introduce: Reasoning Depth, Pattern Recognition, Noise Filtering, and Context Integration provide essential explanatory mechanisms for these variations. Each obfuscation technique targets specific dimensional capabilities: bogus control flow tests reasoning depth through invariant conditions; instruction substitution challenges pattern recognition through equivalent operations; control flow flattening disrupts context integration by fragmenting related code; and combined techniques simultaneously undermine all dimensions. This mapping conclusively explains the inconsistent performance across techniques and establishes a theoretical foundation that advances our understanding of LLM deobfuscation capabilities.
Our taxonomy of five types of error, predicate misinterpretation, structural mapping, control flow structure misinterpretation, arithmetic transformation, and constant propagation errors, provides a comprehensive classification system to analyze LLM limitations. Each error pattern reveals dimensional weaknesses, from the inability to recognize mathematical invariants to the fabrication of constants absent from the original code. This taxonomy represents a significant contribution to understanding the fundamental challenges in machine-driven code analysis.
We establish a three-tier resistance model categorizing techniques according to their effectiveness against LLM-based analysis. Low-resistance techniques challenge single dimensions where advanced models excel; moderate-resistance techniques require integrated capabilities across multiple dimensions; and high-resistance techniques simultaneously challenge all dimensions, currently exceeding the autonomous capabilities of state-of-the-art models. This model provides empirically-grounded insights for understanding protection mechanism effectiveness.
From a security perspective, our findings show that while LLMs reduce expertise barriers for certain reverse engineering tasks, sophisticated protection mechanisms effectively resist fully automated analysis. The universal failure against combined obfuscation techniques defines an upper bound on current capabilities that organizations can exploit when protecting sensitive intellectual property.
The methodological framework established in this investigation, from attacker knowledge levels to dimensional capability mapping, creates a foundation for evaluating emerging capabilities. Future research will expand testing across diverse codebases and obfuscation implementations, develop quantitative metrics for each dimensional capability, and explore hybrid approaches that leverage the complementary strengths of LLMs and traditional program analysis tools.
This work marks a pivotal advancement in the evolution of AI-assisted reverse engineering. Rather than complete automation, these findings point to a new paradigm of human-AI collaboration for complex deobfuscation workflows. The dimensional framework established here enables security practitioners to implement more effective strategies for both code protection and analysis in an increasingly AI-augmented cybersecurity landscape.